# SlimDiff: Training-Free, Activation-Guided Hands-free Slimming of Diffusion Models

### v1

@@ 2.4 Module-Aligned Data-Aware Compression (MADAC)
A core challenge in DM compression is that conventional per-matrix factorization treats weights in isolation, overlooking the functional coupling across attention and feed-forward modules. MADAC addresses this by treating each module as an integrated unit and learning a joint, data-aware decomposition that preserves the end-to-end mapping under empirically observed activation distributions.
@@ 2.5 Automatic Rank Allocation Engine
Given a parameter budget $B$, we choose per–block ranks $\{r_{\ell}\}$ to maximize fidelity with total parameters $\leq B$. Since estimating block-wise utility is infeasible, we use the spectral influence $\mathcal{I}_{\ell,t}$ (Sec. 2.3) as a surrogate and allocate capacity $\propto\sum_{t}\mathcal{I}_{\ell,t}$, so higher-influence blocks retain more rank. Formal mapping details are in App. Sec. A.2.
Softmax-style Allocation. We convert influence scores into retention fractions via a temperature-controlled softmax (App.  A.2), which normalizes importance across layers and concentrates capacity on influential blocks while keeping allocations smooth. Intuitively, each block’s retained rank depends not only on its own importance but also on its relative standing among all blocks, yielding a propagation-aware allocation under the global budget.
@@ 3 Evaluation and Analyses
LD–Pruner (Zhang et al. (2024b)) $\times$ 0.71B 12.37 35.77 0.289 0.22
@@
Table 1 presents quantitative results on generation quality. SlimDiff achieves competitive performance with only minimal degradation: FID of $13.12$ (vs. $13.07$ for SD v1.5) and CLIP score of $0.319$ (vs. $0.322$ for SD v1.5). Notably, SlimDiff is the only method that achieves this performance while being completely backpropagation-free (BP-free), requiring only $500$ data points and 44 A100100 days compared to $6250$ days for the original model. This represents a significant practical advantage over competing methods that require retraining, such as BK-SDM-Base ($13$ A100100 days) and Small Stable Diffusion (extensive retraining on 229229M samples). Note that, for LD-Pruner, results are taken from the paper since the code is not open-sourced, and may not be directly comparable. Figure 2 demonstrates SlimDiff’s ability to preserve semantic content and visual quality across diverse prompts. For the “cute Shiba Inu in a cabbage” prompt, SlimDiff maintains the key semantic elements: the dog’s distinctive features and the cabbage setting, while achieving $27\%$ parameter reduction (760760M vs. 1.041.04B). Artistic prompts like “Van Gogh Starry Night” maintain characteristic brushstrokes and color palettes, while samples for prompts “man staring ahead” and “astronaut bears” highlight robustness across photorealistic and creative domains. More results in App. A.7
@@ 4 Ablation
In this section, we ask: How do design choices affect performance? We ablate SlimDiff’s core components to identify which factors drive quality and efficiency. Specifically, we study (i) SlimSet size, (ii) weighting strategies for timestep correlations, and (iii) the impact of compressing different module types. These analyses clarify why SlimDiff works and where its efficiency gains arise.
@@
(ii) Weighting Strategies for Timestep-Aware Correlation Accumulation. Not all timesteps contribute equally to perceptual fidelity. To evaluate our weighting design for accumulating correlations $\mathcal{C}_{l,t}\rightarrow\mathcal{C}_{l}$, we test four alternatives: (i) uniform averaging across timesteps, (ii) input-activation diversity only, (iii) spectral influence weighting (ours), and (iv) a combined scheme. As shown in Table 6, uniform weighting performs worst, while spectral influence yields the best FID ($13.12$). The combined scheme slightly improves over diversity-only but still falls short of spectral influence. These results highlight that fidelity-aware weighting, captured by the spectral influence score, is essential for effective timestep-aware accumulation.
@@ A.1 Proofs
What is a Column Selection Matrix? A $k$-column selection matrix $\mathrm{M}_{k}\in\{0,1\}^{d\times k}$ has exactly one nonzero per column, indicating the selected indices. For any $A\in\mathbb{R}^{m\times d}$, the product $A\mathrm{M}_{k}$ consists of the $k$ selected columns of $A$. ( Lin et al. (2024); Wang & Singh (2018)). Note that, for any input matrix $X$ and any column selection matrix $\mathrm{M}$, the nonlinearity $\sigma$ satisfies $\sigma(X)\,\mathrm{M}\;=\;\sigma(X\mathrm{M})$. In other words, column selection commutes with $\sigma$. This assumption holds for all activation functions that act elementwise ( Pourkamali-Anaraki & Becker (2019); Gittens & Mahoney (2016)).
@@
where $E_{\text{Nys}}(\mathcal{K})$ denotes the Nyström approximation error ( Gittens & Mahoney (2016) of $\mathcal{K}$ using the same column selection $\mathrm{M}_{k}$.
While our derivation in Eq. 8 holds for any column selection strategy, in practice we adopt CPQR to construct $\mathrm{S}_{k}$. Strong rank-revealing QR Gu & Eisenstat (1996) ensures that the selected columns span a well-conditioned rank-$k$ subspace of $\mathcal{K}$, with Nyström reconstruction error bounded by a modest multiple of the optimal low-rank approximation error Gittens & Mahoney (2016).
@@
Theorem 3 Proof Details: For input $\mathcal{X}$, the $\mathcal{VO}$ interaction is $f(\mathcal{X};\mathcal{W}_{v},\mathcal{W}_{o})=\mathcal{X}\mathcal{W}_{v}\mathcal{W}_{o}$. By whitening with $\mathcal{C}^{1/2}$, the problem becomes SVD of $\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}$. Truncating to rank $r$ yields $\widehat{\mathcal{W}}_{v}=\mathcal{C}^{-1/2}\mathcal{U}_{r},\quad\widehat{\mathcal{W}}_{o}=\Sigma_{r}\mathcal{V}_{r}^{\top}$, with reconstruction error $\sum_{i=r+1}\sigma_{i}^{2}(\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o})$.
We solve our Eq  3 for VO as follows:
$\min_{\widehat{\mathcal{W}}_{v},\,\widehat{\mathcal{W}}_{o}}$ $\sum_{i=1}^{N}\big\|\,f(\mathcal{X}_{i};\mathcal{W}_{v},\mathcal{W}_{o})\;-\;f(\mathcal{X}_{i};\widehat{\mathcal{W}}_{v},\widehat{\mathcal{W}}_{o})\,\big\|_{F}^{2}$ (16)
$=\min_{\widehat{\mathcal{W}}_{v},\,\widehat{\mathcal{W}}_{o}}\sum_{i=1}^{N}\Big\|\,\mathcal{X}_{i}(\mathcal{W}_{v}\mathcal{W}_{o}-\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o})\,\Big\|_{F}^{2}$
$=\min_{\widehat{\mathcal{W}}_{v},\,\widehat{\mathcal{W}}_{o}}\sum_{i=1}^{N}\operatorname{Tr}\!\Big((\mathcal{W}_{v}\mathcal{W}_{o}-\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o})^{\top}\mathcal{X}_{i}^{\top}\mathcal{X}_{i}(\mathcal{W}_{v}\mathcal{W}_{o}-\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o})\Big)$
$=\min_{\widehat{\mathcal{W}}_{v},\,\widehat{\mathcal{W}}_{o}}\Big\|\;\underbrace{\Big(\sum_{i=1}^{N}\mathcal{X}_{i}^{\top}\mathcal{X}_{i}\Big)^{1/2}}_{\mathcal{C}^{1/2}}(\mathcal{W}_{v}\mathcal{W}_{o}-\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o})\;\Big\|_{F}^{2}$
$=\min_{\operatorname{rank}\leq r}\Big\|\;\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}-\mathcal{C}^{1/2}\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o}\;\Big\|_{F}^{2}.$
Optimal Value, Output Matrices: By the Eckart–Young–Mirsky theorem, the optimal solution is given by truncating the SVD of $\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}$ to rank $r$. This yields
$\widehat{\mathcal{W}}_{v}=\mathcal{C}^{-1/2}U_{r},\qquad\widehat{\mathcal{W}}_{o}=\Sigma_{r}V_{r}^{\top},$ (17)
with reconstruction error
$\sum_{i=r+1}\sigma_{i}^{2}(\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}).$ (18)
@@ A.2 Automatic Rank Allocation Engine (Extended)
where $\bar{\phi}\in[0,1]$ is the target average sparsity (determined by budget $B$) and $\varepsilon>0$ is a temperature parameter. The linear term $s_{\ell}\phi_{\ell}$ penalizes sparsifying high-importance blocks, while the entropy regularizer $\phi_{\ell}\log\phi_{\ell}$ prevents winner-take-all collapse by encouraging smooth allocation.
@@
This exponential weighting automatically concentrates capacity on blocks with high TRQ scores while maintaining smooth allocation controlled by temperature $\varepsilon$.
@@
Properties and Guarantees. Our allocation is convex (unique global optimum), propagation-aware (high-influence blocks retain more capacity), and budget-exact (bisection ensures precise cost targeting). Unlike heuristic approaches, the entropy regularization provides principled smoothness while TRQ scores enable cross-family comparison without manual rescaling.
@@ A.3 Related Work
We enumerate related works in this subsection, describing the popular structural pruning methods, alternative methods like reduction of timestep sampling or dynamic token pruning. We also show works on LLMs which use data-aware compression in their pipeline.
@@
Several methods reduce the parameter footprint of diffusion models through pruning or architectural redesign. Block- and layer-pruning approaches compress the U-Net backbone but typically rely on distillation or finetuning to recover fidelity ( Kim et al. (2024); Zhang et al. (2024b, a)). Complementary efforts prune at the timestep or module granularity ( Fang et al. (2023); Yao et al. (2024a)), highlighting the role of sequential error propagation. Our work differs by providing a closed-form, training-free pipeline that operates over functional groups (QK, VO, FFN) with module-aligned objectives.
@@
A separate thread accelerates sampling without changing model size. Attention-driven step reduction ( Wang et al. (2024)) modulates compute over timesteps, while token and cache pruning seek runtime savings ( Bolya & Hoffman (2023)). These methods improve wall-clock latency but add per-run heuristics and do not permanently reduce parameters. SlimDiff is orthogonal: it permanently shrinks dimensions/ranks and can be combined with such inference-time techniques.
@@
Beyond diffusion, activation-aligned and module-aware compression in large transformers ( Lin et al. (2024); Ashkboos et al. (2024)) shows that respecting activation geometry and functional coupling outperforms naive matrix-wise pruning. SlimDiff adapts this perspective to diffusion’s evolving activations: it models per-timestep correlations, weights them by spectral influence, and applies whitening–SVD (QK/VO) and Nyström FFN reductions under a global rank allocator.
@@ A.5 Spectral Influence Score and Diversity Results across layers and modules
We measure the pruning sensitivity of each block using the trace-normalized Rayleigh quotient (TRQ) influence score. TRQ evaluates how strongly a block’s weight operator aligns with the dominant eigenspaces of its activation covariance. Intuitively, it quantifies how effectively a block exploits the signal-rich directions of its input. Because TRQ is normalized by the input trace, the score is directly comparable across blocks of different widths and even across module families ($\mathcal{QK}$, $\mathcal{VO}$, $\mathcal{FFN}$).
Figure 3 highlights several consistent trends. Across all module types, the widest layers (1280 channels) tend to have the lowest importance. For $\mathcal{QK}$ blocks, decoder layers dominate encoder layers, reflecting their exposure to greater variance. In contrast, $\mathcal{FFN}$ blocks show higher importance in the encoder, especially at early timesteps. For $\mathcal{VO}$, cross-attention blocks exhibit stable importance across timesteps, since they depend only on text-side correlations. These observations guide our pruning strategy, where rank allocation is driven directly by TRQ.
We also examine diversity, defined as the spectral spread of input activations. Diversity is highest at early timesteps and decreases as denoising progresses, with mid-U-Net layers achieving higher diversity in fewer steps. This agrees with prior findings that greater diversity reflects less noisy, more semantically aligned activations Wang et al. (2024). However, diversity alone does not reliably predict pruning sensitivity. As confirmed in Table 6, TRQ provides a stronger and more actionable signal for structural pruning, while diversity remains useful primarily as a diagnostic measure.

### v2

@@ 2.4 Module-Aligned Data-Aware Compression (MADAC)
A core challenge in DM compression is that conventional per-matrix factorization treats weights in isolation, overlooking the functional coupling across attention and feed-forward modules. MADAC addresses this by treating each module as an integrated unit and learning a joint, data-aware decomposition that preserves the end-to-end mapping under empirically observed activation distributions. Building on the principles of Lin et al. (2024), we derive diffusion-specific formulations that adapt joint low-rank decomposition to the unique activation geometry of diffusion models.
@@ 2.5 Automatic Rank Allocation Engine
Given a parameter budget $B$, we choose per–block ranks $\{r_{\ell}\}$ to maximize fidelity with total parameters $\leq B$. Since estimating block-wise utility is infeasible, we use the spectral influence $\mathcal{I}_{\ell,t}$ (Sec. 2.3) as a surrogate and allocate capacity $\propto\sum_{t}\mathcal{I}_{\ell,t}$, so higher-influence blocks (layerwise and across denoising trajectory) retain more rank. We first form a timestep mixture $\bar{\mathcal{J}}_{\ell}\simeq\sum_{t}\alpha_{t}\,\mathcal{I}_{\ell,t}$ where $\{\alpha_{t}\}$ are weights that emphasize early denoising steps, reflecting the fact that errors introduced early propagate and amplify downstream. Thus, $\bar{\mathcal{J}}_{\ell}$ is both module-aware and propagation-aware. More details are in App. Sec. A.2.
Softmax-style Allocation. We convert influence scores into retention fractions via a temperature-controlled softmax (App.  A.2), which normalizes importance across layers and concentrates capacity on influential blocks while keeping allocations smooth. Intuitively, each block’s retained rank depends not only on its own importance but also on its relative standing among all blocks and timesteps, yielding a propagation-aware allocation under the global budget.
@@ 3 Evaluation and Analyses
LD–Pruner (Zhang and others (2024)) $\times$ 0.71B 12.37 35.77 0.289 0.22
@@
Table 1 presents quantitative results on generation quality. SlimDiff achieves competitive performance with only minimal degradation: FID of $13.12$ (vs. $13.07$ for SD v1.5) and CLIP score of $0.319$ (vs. $0.322$ for SD v1.5). Notably, SlimDiff is the only method that achieves this performance while being completely backpropagation-free (BP-free), requiring only $500$ data points and 44 A100100 days compared to $6250$ days for the original model. This represents a significant practical advantage over competing methods that require retraining, such as BK-SDM-Base ($13$ A100100 days) and Small Stable Diffusion (extensive retraining on 229229M samples). Note that, for LD-Pruner, results are taken from the paper since the code is not open-sourced, and may not be directly comparable. Figure 3 demonstrates SlimDiff’s ability to preserve semantic content and visual quality across diverse prompts. For the “cute Shiba Inu in a cabbage” prompt, SlimDiff maintains the key semantic elements: the dog’s distinctive features and the cabbage setting, while achieving $27\%$ parameter reduction (760760M vs. 1.041.04B). Artistic prompts like “Van Gogh Starry Night” maintain characteristic brushstrokes and color palettes, while samples for prompts “man staring ahead” and “astronaut bears” highlight robustness across photorealistic and creative domains. More results in App. A.8
@@ 4 Ablation
In this section, we ask: How do design choices affect performance? We ablate SlimDiff’s core components to identify which factors drive quality and efficiency. Specifically, we study (i) SlimSet size, (ii) weighting strategies for timestep correlations, and (iii) the impact of compressing different module types. These analyses clarify why SlimDiff works and where its efficiency gains arise. Beyond these core ablations, Appx. 7 analyzes SlimSet coverage across datasets, Appx. 10 details per-block compression strategies, and Appx. 12 reports SlimDiff’s extension to quantized baselines; we collate the main takeaways from these additional studies under (iv) Extended Empirical Studies.
@@
(ii) Weighting Strategies for Timestep-Aware Correlation Accumulation. Not all timesteps contribute equally to perceptual fidelity. To evaluate our weighting design for accumulating correlations $\mathcal{C}_{l,t}\rightarrow\mathcal{C}_{l}$, we test four alternatives: (i) uniform averaging across timesteps, (ii) input-activation diversity only, (iii) spectral influence weighting (ours), and (iv) a combined scheme. As shown in Table 6, uniform weighting performs worst, while spectral influence yields the best FID ($13.12$). The combined scheme slightly improves over diversity-only but still falls short of spectral influence. These results highlight that fidelity-aware weighting, captured by the spectral influence score, is essential for effective timestep-aware accumulation of correlation matrices.
@@
Future Work. SlimDiff is extensible to DiT-style backbones such as Stable Diffusion 3.5’s MMDiT (Appx. 11) via an RMSNorm-aware, data-driven QK decomposition tailored to jointly normalized query–key streams. A complementary direction is to refine propagation-aware rank allocation in joint attention, where image and text share heads but not FFNs, to more carefully balance cross-modal capacity under a single parameter budget.
@@ A.1 Proofs
What is a Column Selection Matrix? A $k$-column selection matrix $\mathrm{M}_{k}\in\{0,1\}^{d\times k}$ has exactly one nonzero per column, indicating the selected indices. For any $A\in\mathbb{R}^{m\times d}$, the product $A\mathrm{M}_{k}$ consists of the $k$ selected columns of $A$. ( Lin et al. (2024); Wang and Singh (2018)). Note that, for any input matrix $X$ and any column selection matrix $\mathrm{M}$, the nonlinearity $\sigma$ satisfies $\sigma(X)\,\mathrm{M}\;=\;\sigma(X\mathrm{M})$. In other words, column selection commutes with $\sigma$. This assumption holds for all activation functions that act elementwise ( Pourkamali-Anaraki and Becker (2019); Gittens and Mahoney (2016)).
@@
where $E_{\text{Nys}}(\mathcal{K})$ denotes the Nyström approximation error ( Gittens and Mahoney (2016) of $\mathcal{K}$ using the same column selection $\mathrm{M}_{k}$.
While our derivation in Eq. 10 holds for any column selection strategy, in practice we adopt CPQR to construct $\mathrm{S}_{k}$. Strong rank-revealing QR Gu and Eisenstat (1996) ensures that the selected columns span a well-conditioned rank-$k$ subspace of $\mathcal{K}$, with Nyström reconstruction error bounded by a modest multiple of the optimal low-rank approximation error Gittens and Mahoney (2016).
@@
Theorem 3 (Adapted from ModeGPT( Lin et al. (2024))): ModeGPT introduced a correlation-aware formulation showing that, under the metric defined by $\mathcal{C}$, the optimal low-rank approximation of the composite $\mathcal{W}_{v}\mathcal{W}_{o}$ is obtained by truncating the SVD of the whitened operator $\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}$.
Diffusion-Specific Adaptation: Extending this principle to diffusion models, we specialize the modular objective in Eq. 5 for the $\mathcal{VO}$ module. Here, the coupling between value and output projections differs across attention types: in self-attention, both $\mathcal{W}_{v}$ and $\mathcal{W}_{o}$ operate on latent features, whereas in cross-attention, $\mathcal{W}_{v}$ originates from text-conditioning embeddings that interact with latent activations through $\mathcal{W}_{o}$. Under this setup, our optimization objective becomes:
$\min_{\operatorname{rank}\leq r}\Big\|\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}-\mathcal{C}^{1/2}\widehat{\mathcal{W}}_{v}\widehat{\mathcal{W}}_{o}\Big\|_{F}^{2},$ (18)
whose optimal solution follows directly from the truncated SVD of $\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}$, yielding
$\widehat{\mathcal{W}}_{v}=\mathcal{C}^{-1/2}U_{r},\qquad\widehat{\mathcal{W}}_{o}=\Sigma_{r}V_{r}^{\top},$
and reconstruction error
$\sum_{i=r+1}\sigma_{i}^{2}(\mathcal{C}^{1/2}\mathcal{W}_{v}\mathcal{W}_{o}).$
@@ A.2 Automatic Rank Allocation Engine (Extended)
where $\bar{\phi}\in[0,1]$ is the target average sparsity (determined by budget $B$) and $\varepsilon>0$ is a temperature parameter. The linear term $s_{\ell}\phi_{\ell}$ penalizes sparsifying high-importance blocks as well as the early denoising steps (error accumulation-aware), while the entropy regularizer $\phi_{\ell}\log\phi_{\ell}$ prevents winner-take-all collapse by encouraging smooth allocation.
@@
This exponential weighting automatically concentrates capacity on blocks with high TRQ scores and early denoising step-aware while maintaining smooth allocation controlled by temperature $\varepsilon$.
@@
Properties and Guarantees. Our allocation is convex (unique global optimum), propagation-aware (high-influence blocks (in U-Net structure as well as in early denoising step) retain more capacity), and budget-exact (bisection ensures precise cost targeting). Unlike heuristic approaches, the entropy regularization provides principled smoothness while TRQ scores enable cross-family comparison without manual rescaling.
@@ A.3 Related Work
We enumerate related works in this subsection, describing the popular structural pruning/factorization methods, alternative methods like reduction of timestep sampling or dynamic token pruning. We also show works on LLMs which use data-aware compression in their pipeline.
@@
Several methods reduce the parameter footprint of diffusion models through pruning or architectural redesign. Block- and layer-pruning approaches compress the U-Net backbone but typically rely on distillation or finetuning to recover fidelity ( Kim et al. (2024); Zhang and others (2024); Zhang et al. (2024)). Complementary efforts prune at the timestep or module granularity ( Fang et al. (2023); Yao et al. (2024)), highlighting the role of sequential error propagation. Our work differs by providing a closed-form, training-free pipeline that operates over functional groups (QK, VO, FFN) with module-aligned objectives.
@@
A separate thread accelerates sampling without changing model size. Attention-driven step reduction ( Wang et al. (2024a)) modulates compute over timesteps, while token and cache pruning seek runtime savings ( Bolya and Hoffman (2023)). These methods improve wall-clock latency but add per-run heuristics and do not permanently reduce parameters. SlimDiff is orthogonal: it permanently shrinks dimensions/ranks and can be combined with such inference-time techniques.
@@
Beyond diffusion, activation-aligned and module-aware compression in large transformers ( Lin et al. (2024); Ashkboos et al. (2024)) shows that respecting activation geometry and functional coupling outperforms naive matrix-wise dimension reduction. SlimDiff adapts this perspective to diffusion’s evolving activations: it models per-timestep correlations, weights them by spectral influence, and applies whitening–SVD (QK/VO) and Nyström FFN reductions under a global rank allocator.
@@ A.5 Spectral Influence Score and Diversity Results across layers and modules
We measure the compression sensitivity of each block using the trace-normalized Rayleigh quotient (TRQ) influence score. TRQ evaluates how strongly a block’s weight operator aligns with the dominant eigenspaces of its activation covariance. Intuitively, it quantifies how effectively a block exploits the signal-rich directions of its input. Because TRQ is normalized by the input trace, the score is directly comparable across blocks of different widths and even across module families ($\mathcal{QK}$, $\mathcal{VO}$, $\mathcal{FFN}$).
Figure 4 highlights several consistent trends. Across all module types, the widest layers (1280 channels) tend to have the lowest importance. For $\mathcal{QK}$ blocks, decoder layers dominate encoder layers, reflecting their exposure to greater variance. In contrast, $\mathcal{FFN}$ blocks show higher importance in the encoder, especially at early timesteps. For $\mathcal{VO}$, cross-attention blocks exhibit stable importance across timesteps, since they depend only on text-side correlations. These observations guide our compression strategy, where rank allocation is driven directly by TRQ.
We also examine diversity, defined as the spectral spread of input activations. Diversity is highest at early timesteps and decreases as denoising progresses, with mid-U-Net layers achieving higher diversity in fewer steps. This agrees with prior findings that greater diversity reflects less noisy, more semantically aligned activations Wang et al. (2024a). However, diversity alone does not reliably predict compression sensitivity. As confirmed in Table 6, TRQ provides a stronger and more actionable signal for structural rank reduction, while diversity remains useful primarily as a diagnostic measure.